OpenStudy (anonymous):

Simplify 5k(2k3 - 4k2 + 4k - 3)

OpenStudy (unklerhaukus):

\[5k\big(2k^3 - 4k^2 + 4k - 3\big)\\ =5k\times(2k^3)-5k\times(4k^2)+5k\times(4k)-5k\times(3)\\=\]